X(0) -90020 4) Solve for lups - [-3x K te 42 12x7 at
Answer #1

dy dx = 1-321 2+2x-y dt at dy A hp XP 9+27- 1-32 ar al hp xp 2+2x-y . 1-3x dy 2 (17) - = da 1-30 1-34 dy da 5 2 (173) 1-37 1-H 기 da y. (1-333 (-3% -3기스 43 -3d1=3+?ad 3H y +7 ts - 8 3 ata & CKE-2) ht to hts to O se 들 CE 드 J- 금에 3 3 성 2. 0 -2+) f 2라 4